A manufacturer is interested in the output voltage of a power supply used in a PC. Output voltage is assumed to be normally distributed, with standard deviation 0.36 Volts, and the manufacturer wishes to test H0: µ = 5 Volts against H1: µ ≠ 5 Volts, using n = 6 units. a-The acceptance region is 4.85 ≤ x-bar ≤ 5.12. Find the value of α. b-Find the power of the test for detecting a true mean output voltage of 5.1 Volts
Answer #1

a)

value of α =1-P(4.85<X<5.12)=1-P((4.85-5)*sqrt(6)/0.36<Z<(5.12-5)*sqrt(6)/0.36)

=1-P(-1.02<Z<0.82)=1-(0.7939-0.1539)=0.3600

b)

power of the test:=1-P(4.85<X<5.12)=1-P((4.85-5.1)*sqrt(6)/0.36<Z<(5.12-5.1)*sqrt(6)/0.36)

=1-P(-1.70<Z<0.14)=1-(0.5557-0.0446)=0.4889